OpenStudy (anonymous):

explain why f(x)=f^-1(x) for any function of the form f(x)=-x+b

OpenStudy (e.mccormick):

\( f(x)=f^{-1}(x)\) for any function of the form f(x)=-x+b Did you put y in and solve and see what it does?

OpenStudy (anonymous):

this has to do with the inverse : so lets say y=2 and b=5 x=-2+5 x=3 is that what u want ?

OpenStudy (e.mccormick):

No, the method for finding the inverse. Swap x and y, then solve for y.

OpenStudy (e.mccormick):

The steps you do when you find the inverse will show why this is true.

OpenStudy (anonymous):

when we have 2=-x+5 then we swap x and y we get the same answer as y= 3+5 , i still dont get why is there ^-1 beside f

OpenStudy (anonymous):

thx anyways :) i think i got it and my question was dumb .

OpenStudy (e.mccormick):

Hehe. The dumb question is the one you do not ask. f(x)=-x+b Swap x/y to find inverse. x=-y+b x-b=-y+b-b x-b=-y (-1)(x-b)=(-1)-y -x+b=y therefore f inverse is: y=-x+b. Explaining that process answers the question.

OpenStudy (anonymous):

Thx :)

OpenStudy (e.mccormick):

np. have fun!